OpenStudy (anonymous):

simplify the number using the imaginary unit i √(-25)

zepdrix (zepdrix):

Recall that we define the imaginary unit this way: \(\large\rm \color{orangered}{\sqrt{-1}=i}\)

zepdrix (zepdrix):

So let's break down our sqrt(-25) and try to find a sqrt(-1) in it.

zepdrix (zepdrix):

\[\large\rm \sqrt{-25}=\sqrt{-1\cdot25}=\sqrt{-1}\cdot\sqrt{25}\]Do you understand those steps? :o

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

@zepdrix once you get that what do you do next?

zepdrix (zepdrix):

\[\large\rm \sqrt{-25}=\sqrt{-1\cdot25}=\color{orangered}{\sqrt{-1}}\cdot\sqrt{25}\]Well, we were able to find a sqrt(-1) in this mess, which we defined to be this imaginary unit i,\[\large\rm =\color{orangered}{i}\cdot\sqrt{25}\]Then proceed further by simplifying the square root of 25.

OpenStudy (anonymous):

@zepdrix so itd be 5i?

zepdrix (zepdrix):

yes c: